# Benchmarking Knowledge Boundary for Large Language Model: A Different Perspective on Model Evaluation

### v1

@@ Abstract
In recent years, substantial advancements have been made in the development of large language models, achieving remarkable performance across diverse tasks. To evaluate the knowledge ability of language models, previous studies have proposed lots of benchmarks based on question-answering pairs. We argue that it is not reliable and comprehensive to evaluate language models with a fixed question or limited paraphrases as the query, since language models are sensitive to prompt. Therefore, we introduce a novel concept named knowledge boundary to encompass both prompt-agnostic and prompt-sensitive knowledge within language models. Knowledge boundary avoids prompt sensitivity in language model evaluations, rendering them more dependable and robust. To explore the knowledge boundary for a given model, we propose projected gradient descent method with semantic constraints, a new algorithm designed to identify the optimal prompt for each piece of knowledge. Experiments demonstrate a superior performance of our algorithm in computing the knowledge boundary compared to existing methods. Furthermore, we evaluate the ability of multiple language models in several domains with knowledge boundary.
@@ 1 Introduction
As shown in Figure 1, from the perspective of the model’s mastery of the textual form of knowledge, knowledge can be divided into three classes: 1) Prompt-Agnostic Knowledge that can be correctly answered for any textual form; 2) Prompt-Sensitive Knowledge that is sensitive to the form of the prompt fed into the model; 3) Unanswerable Knowledge that is unable to be answered by the model, regardless of the prompt employed. The majority of previous research on model evaluation ignored the presence of Prompt-Sensitive Knowledge, resorting to oversimplified binary evaluations, classifying the model’s knowledge mastery merely as true or false. Dong et al. (2023) attempts to assess LLM through diverse paraphrases, yet these evaluations remain confined to limited textual forms of knowledge.
In this paper, we aim to reduce the contingency when evaluating LLMs. Different from previous paradigms of LLM evaluation, we attempt to explore these Unanswerable Knowledge of the model to be evaluated, thereby illuminating the knowledge boundaries of LLMs. How can we find Unanswerable knowledge for the model? It is obvious that trying all prompts for the knowledge to query the model is too resource-intensive. Therefore, we choose to make efforts to search for the optimal prompt. We formalize optimal prompt searching as a discrete optimization problem: given some question paraphrases, we search for a prompt to maximize the probability of generating the correct answer. We propose Projected Gradient Descent method with Constraints (PGDC), a new algorithm that updates prompt with gradient descent and implements proximal projection to search discrete prompts. To ensure that the optimized prompt has the same semantics as the original prompt, we introduce the semantic loss, which is a measure of the distance between the semantic representations of the optimized prompt and the original prompt.
@@
In summary, our contributions are: (1) We propose a new evaluation paradigm for benchmarking knowledge boundaries to compare models’ capabilities, which can reduce the randomness in current evaluations. (2) We design PGDC, projected gradient descent method with constraints, to optimize prompts and obtain knowledge boundaries of LLMs which achieves the best results on four datasets. (3) We evaluate five models using knowledge boundaries and obtain some valuable findings.
Our code and data will be released to the community to facilitate future research.
@@ 3 Method
To obtain a better knowledge boundary for LLM, our effort is directed towards identifying the optimal prompt within the semantic space. As illustrated in Figure 2, PGDC optimizes prompt in the neighbour semantic space of the original question. The prompt in text form is first mapped to prompt in embedding form as continuous text embedding. During PGDC optimization, the text embedding is updated through gradient descent with direction of the loss function $\Phi(\cdot)$. After each update, if the text embedding is close to a discrete prompt, it is projected to the discrete prompt through embedding projection. To avoid the text embedding from entering unprojectable region where there are no close discrete prompts to project, we introduce a regularization to force the embedding not to enter these regions. After multiple iterations of update, we get the final optimized prompt.
@@ 3.1 PGDC Optimization
$\centering\Phi(X)=L(X,A)+\lambda_{1}R(X,Q)+\lambda_{2}\delta(X)\@add@centering$ (6)
@@ 4 Experiments
Table 1: The success rate of constructing prompts to elicit specific knowledge on four Datasets. We conduct experiments on four different LLMs to illustrate the performance of our proposed PGDC. Dataset PaRaRel and KAssess provide true knowledge to characterize the ability of different methods on obtaining knowledge boundary while the pieces of knowledge in dataset CFACT and AlCuna are fake which shows the robustness of PGDC.
@@ 4.1 Datasets and Models
To test whether our optimized prompts leak answers or induce hallucinations that cause LLMs to answer knowledge that they could not answer originally, we perform evaluations on two counterfactual datasets, COUNTERFACT (denoted as CFACT) (Meng et al., 2022) and AlCuna (Yin et al., 2023). CFACT containes 20K counterfactual knowledge records with a diverse set of subjects, relations, and linguistic variations. AlCuna is a biological dataset for evaluating the ability of the model in face of new knowledge.
@@ 4.4 Comparison with Prompt Optimization Method
As shown in Table 3, we can find that Autoprompt induces the model to output target answers on counterfactual datasets in a large percentage. This result suggests that Autoprompt is more similar to an adiversarial attack algorithm that are committed to getting the target answer, while PGDC optimizes the prompt within the semantic constraint.
@@ 5.2 Results
We can find that Mistral has the largest knowledge boundaries overall. LLaMA2 exceeds the other models by a lot in engineering domain. It may be because LLaMA2 uses a lot more new labeled code data for training. Vicuna performs not far behind LLaMA2 on other topics. GPT-2 has very small knowledge boundaries and performs poorly in the more specialized medical domain. By identifying more reliable knowledge boundaries, we help select the appropriate LLM in practice.
However, it is worth noting that these scores are quite low (around 20 points). This is due to the fact that we use more difficult cloze-style questions for reliability. It also reflects the fact that the results obtained from choice-style benchmark may be too high. The problems in MMLU are relatively specialized, and the current general-purpose models do not have a lot of knowledge in the relevant areas.
@@ A.1 Datasets
AlCuna (Yin et al., 2023) is used for evaluating the ability of LLMs in face of new knowledge which consists of a total of 84351 questions about 3554 artificial entities. We only select the cloze-style portion of the questions to be used for the experiments.
KAssess
Zero-Shot
1. 10,000 metres record is held by [Kenenisa Bekele]
Few-Shots
@@
Discriminator
3. Check whether the following statement is correct: 10,000 metres record is held by Kenenisa Bekele. The statement is (True/False): [True]
PaRaRel
Zero-Shot
1. The mother tongue of Go Hyeon-jeong is [Korean]
Few-Shots
@@ 1. The mother tongue of Go Hyeon-jeong is [Korean]
Discriminator
3. Check whether the following statement is correct: The mother tongue of Go Hyeon-jeong is Korean. The statement is (True/False): [True]
CFACT
Zero-Shot
1. IBM Connections is created by [Adobe]
Few-Shots
@@ 1. IBM Connections is created by [Adobe]
Discriminator
3. Check whether the following statement is correct: IBM Connections is created by Adobe. The statement is (True/False): [True]
AlCuna
Zero-Shot
1. What’s the body length of Leuciaiaivea? [8.1 cm]
Few-Shots
@@
Discriminator
3. Check whether the following statement is correct: What’s the body length of Leuciaiaivea? 8.1 cm. The statement is (True/False): [True]
@@ Appendix F Human Evaluation
Support and Reference: If you encounter any confusion regarding professional knowledge or context while performing this task, please feel free to reach out to us for clarification.You may also refer to Wikipedia or other reliable sources to gain further understanding.
Feedback Mechanism: You can directly submit your queries, concerns, or suggestions to us.

### v2

# Benchmarking Knowledge Boundary for Large Language Models: A Different Perspective on Model Evaluation

@@ Abstract
In recent years, substantial advancements have been made in the development of large language models, achieving remarkable performance across diverse tasks. To evaluate the knowledge ability of language models, previous studies have proposed lots of benchmarks based on question-answering pairs. We argue that it is not reliable and comprehensive to evaluate language models with a fixed question or limited paraphrases as the query, since language models are sensitive to prompt. Therefore, we introduce a novel concept named knowledge boundary to encompass both prompt-agnostic and prompt-sensitive knowledge within language models. Knowledge boundary avoids prompt sensitivity in language model evaluations, rendering them more dependable and robust. To explore the knowledge boundary for a given model, we propose a projected gradient descent method with semantic constraints, a new algorithm designed to identify the optimal prompt for each piece of knowledge. Experiments demonstrate a superior performance of our algorithm in computing the knowledge boundary compared to existing methods. Furthermore, we evaluate the ability of multiple language models in several domains with knowledge boundary.
@@ 1 Introduction
As shown in Figure 1, from the perspective of the model’s mastery of the textual form of knowledge, knowledge can be divided into three classes: 1) Prompt-Agnostic Knowledge that can be correctly answered for any textual form; 2) Prompt-Sensitive Knowledge that is sensitive to the form of the prompt fed into the model; 3) Unanswerable Knowledge that is unable to be answered by the model, regardless of the prompt employed. The majority of previous research on model evaluation ignored the presence of Prompt-Sensitive Knowledge, resorting to oversimplified binary evaluations, classifying the model’s knowledge mastery merely as true or false. Dong et al. (2023) attempts to assess LLM through diverse paraphrases, yet these evaluations remain confined to limited textual forms of knowledge. We give strict definitions of three types of knowledge in Section 2.1.
In this paper, we aim to reduce the contingency when evaluating LLMs. Different from previous paradigms of LLM evaluation, we attempt to explore the Unanswerable Knowledge of the model to be evaluated, thereby illuminating the knowledge boundaries of LLMs. How can we find Unanswerable knowledge for the model? It is obvious that trying all prompts for the knowledge to query the model is too resource-intensive. Therefore, we choose to make efforts to search for the optimal prompt. We formalize optimal prompt searching as a discrete optimization problem: given some question paraphrases, we search for a prompt to maximize the probability of generating the correct answer. We propose the Projected Gradient Descent method with Constraints (PGDC), a new algorithm that updates prompt with gradient descent and implements proximal projection to search discrete prompts. To ensure that the optimized prompt has the same semantics as the original prompt, we introduce semantic loss, which is a measure of the distance between the semantic representations of the optimized prompt and the original prompt.
@@
In summary, our contributions are: (1) We propose a new evaluation paradigm for benchmarking knowledge boundaries to compare models’ capabilities, which can reduce the randomness in current evaluations. (2) We design PGDC, a projected gradient descent method with constraints, to optimize prompts and obtain knowledge boundaries of LLMs which achieves the best results on four datasets. (3) We evaluate five models using knowledge boundaries and obtain some valuable findings.
Our code and data are released to facilitate future research [^1] https://github.com/pkulcwmzx/knowledge boundary.
@@ 2 Preliminaries
2.1 Knowledge Boundary Definition
First, we provide a strict definition of the three types of knowledge in LLMs, Prompt-Agnostic Knowledge, Prompt-Sensitive Knowledge and Unanswerable Knowledge.
Formally, let $k$ denote a piece of knowledge, and let $\Theta$ represent an LM. We assess whether $\Theta$possesses" knowledge of k by calculating $P(a_{k}^{i}|q_{k}^{i},\Theta)$, where $q_{k}^{i}\in Q_{k},a_{k}^{i}\in A_{k}$ represents a pair of input-output tokens to verify the knowledge $k$.
For the universe of all conceivable knowledge $U$, and a threshold $\epsilon$ within the range $(0.5,1]$, the three types of knowledge for model $\Theta$ are defined as follows:
  • Prompt-Agnostic Knowledge: $K^{A}=\{k\in U|P(a_{k}^{i}|q_{k}^{i},\Theta)>\epsilon,\forall q_{k}^{i}\in Q_{k},\forall a_{k}^{i}\in A_{k}\}$
  • Unanswerable Knowledge: $K^{U}=\{k\in U|P(a_{k}^{i}|q_{k}^{i},\Theta)<\epsilon,\forall q_{k}^{i}\in Q_{k},\forall a_{k}^{i}\in A_{k}\}$
  • Prompt-Sensitive Knowledge: $K^{S}=\{k\in U|k\notin K^{U}\land k\notin K^{A}\}$
In short, for a piece of knowledge $k$, if the model $\Theta$ is able to answer the question about $k$ correctly under at least one expression, $k$ is inside the knowledge boundary of $\Theta$. If $\Theta$ is unable to answer the question about $k$ correctly under any expressions, $k$ is outside the knowledge boundary of the model.
@@ 3 Method
To obtain a better knowledge boundary for LLM, our effort is directed toward identifying the optimal prompt within the semantic space. As illustrated in Figure 2, PGDC optimizes prompt in the neighbor semantic space of the original question. The prompt in text form is first mapped to prompt in embedding form as continuous text embedding. During PGDC optimization, the text embedding is updated through gradient descent with direction of the loss function $\Phi(\cdot)$. After each update, if the text embedding is close to a discrete prompt, it is projected to the discrete prompt through embedding projection. To avoid the text embedding from entering unprojectable region where there are no close discrete prompts to project, we introduce a regularization to force the embedding not to enter these regions. After multiple iterations of the update, we get the final optimized prompt.
@@ 3.1 PGDC Optimization
$\centering\Phi(X)=L(X,A)+\lambda_{1}R(X,Q)+\lambda_{2}\delta(X).\@add@centering$ (6)
@@ 4 Experiments
Table 1: The success rate of constructing prompts to elicit specific knowledge on four Datasets. We conduct experiments on four different LLMs to illustrate the performance of our proposed PGDC. Dataset PaRaRel and KAssess provide true knowledge to characterize the ability of different methods to obtain knowledge boundary while the pieces of knowledge in dataset CFACT and AlCuna are fake which shows the robustness of PGDC.
@@ 4.1 Datasets and Models
To test whether our optimized prompts leak answers or induce hallucinations that cause LLMs to answer knowledge that they could not answer originally, we perform evaluations on two counterfactual datasets, COUNTERFACT (denoted as CFACT) (Meng et al., 2022) and AlCuna (Yin et al., 2023). CFACT contains 20K counterfactual knowledge records with a diverse set of subjects, relations, and linguistic variations. AlCuna is a biological dataset for evaluating the ability of the model in face of new knowledge.
@@ 4.4 Comparison with Prompt Optimization Method
As shown in Table 3, we can find that Autoprompt induces the model to output target answers on counterfactual datasets in a large percentage. This result suggests that Autoprompt is more similar to an adversarial attack algorithm that is committed to getting the target answer, while PGDC optimizes the prompt within the semantic constraint.
@@ 5.2 Results
We can find that Mistral has the largest knowledge boundaries overall. LLaMA2 exceeds the other models by a lot in the engineering domain. It may be because LLaMA2 uses a lot more new labeled code data for training. Vicuna performs not far behind LLaMA2 on other topics. GPT-2 has very small knowledge boundaries and performs poorly in the more specialized medical domain. By identifying more reliable knowledge boundaries, we help select the appropriate LLM in practice.
However, it is worth noting that these scores are quite low (around 20 points). This is due to the fact that we use more difficult cloze-style questions for reliability. It also reflects the fact that the results obtained from the choice-style benchmark may be too high. The problems in MMLU are relatively specialized, and the current general-purpose models do not have a lot of knowledge in the relevant areas.
6 Discussion of Randomness in Model Evaluation
The inherent randomness in model evaluation presents significant challenges to the reliability of evaluation results and the process of model selection. This randomness can be seen in several key areas:
Benchmark Question Selection
Randomness is inherent in the selection of benchmark questions for testing, as the objective is to choose representative questions that accurately demonstrate the model’s abilities in specific areas or domains (Guo et al., 2023; Zhong et al., 2023).
Prompt Expression
Formulating the selected questions into prompts introduces additional randomness, including aspects such as the phrasing of questions, system instructions, the order of options, and other details. Numerous studies have demonstrated that models are sensitive to variations in prompts (Ji et al., 2023; Maharana et al., 2023; Chang and Bergen, 2023; Chen et al., 2023).
Decoding Settings
The choice of decoding parameters (e.g., temperature) during evaluation can introduce randomness into the model’s output, potentially raising concerns about fairness.
Output Evaluation
The evaluation of the model’s generated content also involves randomness. Some studies use exact matching (Yin et al., 2023), while others employ GPT-4 for evaluation (Chiang et al., 2023); however, neither method is flawless.
Given the various sources of randomness in model evaluation, developing reliable, stable, and informative evaluation methods remains a significant challenge and warrants further research. To address this concern, we introduce the concept of the "knowledge boundary," which employs optimized prompts to test the upper limits of a model’s capabilities, thereby minimizing the randomness associated with prompt selection.
@@ 8 Conclusion
Additionally, despite our efforts to ensure the truthfulness of optimized prompts, there remains a small probability that the semantics of the prompt may change, especially for weaker language models such as GPT-2.
@@
Acknowledgements
This work was supported by Beijing Science and Technology Program (Z231100007423011), National Key R&D Program of China (2021YFF0901502), National Science Foundation of China (No. 62161160339) and Key Laboratory of Science, Technology and Standard in Press Industry (Key Laboratory of Intelligent Press Media Technology). We would like to thank Zhidong Jia, Yezhen Chen, Baizhou Huang, Junzhe Zhang and members of the group for their valuable feedback and discussions. We appreciate the anonymous reviewers for their helpful comments. Xiaojun Wan is the corresponding author.
@@ A.1 Datasets
AlCuna (Yin et al., 2023) is used for evaluating the ability of LLMs in the face of new knowledge which consists of a total of 84351 questions about 3554 artificial entities. We only select the cloze-style portion of the questions to be used for the experiments.
KAssess Zero-Shot 1. 10,000 metres record is held by [Kenenisa Bekele]
@@
Discriminator 3. Check whether the following statement is correct: 10,000 metres record is held by Kenenisa Bekele. The statement is (True/False): [True]
PaRaRel Zero-Shot 1. The mother tongue of Go Hyeon-jeong is [Korean]
@@
Discriminator 3. Check whether the following statement is correct: The mother tongue of Go Hyeon-jeong is Korean. The statement is (True/False): [True]
CFACT Zero-Shot 1. IBM Connections is created by [Adobe]
@@
Discriminator 3. Check whether the following statement is correct: IBM Connections is created by Adobe. The statement is (True/False): [True]
AlCuna Zero-Shot 1. What’s the body length of Leuciaiaivea? [8.1 cm]
@@
Discriminator 3. Check whether the following statement is correct: What’s the body length of Leuciaiaivea? 8.1 cm. The statement is (True/False): [True]
@@ Appendix F Human Evaluation
Support and Reference: If you encounter any confusion regarding professional knowledge or context while performing this task, please feel free to reach out to us for clarification.You may also refer to Wikipedia or other reliable sources to gain further understanding. Feedback Mechanism: You can directly submit your queries, concerns, or suggestions to us.